OpenStudy (anonymous):

Tangent 7pi/8 using half angle formula

OpenStudy (jdoe0001):

I assume you have your half-angle formulas sheet?

OpenStudy (jdoe0001):

:)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

This one is meant for tangents :) http://oakroadsystems.com/twt/double.htm

OpenStudy (anonymous):

I know how to do it but I can't get the right answer.

OpenStudy (anonymous):

I can't help with that @Jackiedavis Sorry :( I am not taking calculus or precalculus yet @Mertsj and @Luis_Rivera Could u please help @Jackiedavis out? Thanks :)

OpenStudy (anonymous):

Thanks for trying

OpenStudy (anonymous):

Your welcome :(, I will try to message a moderator to help

OpenStudy (jdoe0001):

what answer did you get Jackiedavis?

OpenStudy (anonymous):

2-sqrt 2 all over 2

OpenStudy (jdoe0001):

I get,... one sec

OpenStudy (jdoe0001):

\[tan(\frac{7\pi}{8})= tan(\frac{\frac{7\pi}{4}}{2})=\sqrt(\frac{1-cos(\frac{7\pi}{4})}{1+cos(\frac{7\pi}{4}})\]

OpenStudy (jdoe0001):

and thus I get \[\frac{2-\sqrt(2)}{2+\sqrt(2)}\]

OpenStudy (jdoe0001):

woops, sqrt of that :)